If y = log (1+x/1-x)^1/4-1/2 (tan^-1 x),then dy/DX is equal to

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
y = log(1 + x/(1-x))1/4 – 1/2.tan-1x = 1/4.log(1/(1-x)) – 1/2.tan-1x = 1/4.log(x – 1) – 1/2.tan-1x.
y' = 1/4(x-1) – 1/2(1+x2).
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)